• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name paschke

    Log entry time 18:15:24 on August29,2009

    Entry number 286871

    keyword=target update

    Report from Dave Meekins: cryoflow from CHL reduced. This isn't a fine-tunable process, and will probably drift back up a little. At present level, heater running between 600-800W. There is a plan for getting sufficient heater power overhead on monday (by daisy chaining power supplies), both for the PID loop controller and for the auxilliary supplies in the counting house. We should be able to return to higher cryo flow after this configuration change on monday, with a safe margin of power.

    Problem today was a result of a blown fuse on the power supply downstairs. That power supply is being used inside specs, but did have problems today.